E and F are mutually exclusive events. P(E) = 0.15; P(F) = 0.71. Find P(E | F)

Found 2 solutions by greenestamps, Edwin McCravy:
Answer by greenestamps(13195)



P(E|F) is the probability that E occurs, given that F occurs. If E and F are mutually exclusive, then that probability is 0.

Mathematically,

P(E|F) = P(E and F)/P(F) = 0/0.71 = 0

Answer by Edwin McCravy(20054)

There are two meanings of "exclusive", so when you see the term "mutually
exclusive", don't think of this meaning of "exclusive":

socially restricted (as of the upper class)

Instead think of this meaning of "exclusive":

excluding

Events that are mutually exclusive EXCLUDE EACH OTHER.  That means if one
occurs, the other is EXCLUDED.  That means if one occurs, the other CANNOT
occur.
E and F are mutually exclusive events.

So that tells us that the probability of them both happening is impossible,
which mean P(E and F) = 0

So, this is really a trick question.  It doesn't matter what their probabilities
are, they CANNOT both occur.  P(E|F) asks for the probability that E will occur
if we are given that F is known to occur.  So the answer must be ZERO!!!

Answer: 0

Edwin